The heights of grade 11 male students are normally distributed with a mean of 65 inches and a standard deviation 2.25 inches. (a) Find the values of 45.75 inches and 72.5 inches. (b) What is the probability that a randomly chosen member of the group has height x berween 60 inches and 70 inches?

1
Expert's answer
2022-04-25T10:44:06-0400

We have a normal distribution, "\mu=65, \sigma=2.25."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"\text{(a) }z_1=\cfrac{45.75-65}{2.25}=-8.56;\\
z_2=\cfrac{72.5-65}{2.25}=3.33.\\
\text{(b) }z_3=\cfrac{60-65}{2.25}=-2.22;\\
z_4=\cfrac{70-65}{2.25}=2.22.\\
P(60<X<70)=\\
=P(-2.22<Z<2.22)=\\
=P(Z<2.22)-P(Z<-2.22)=\\
=0.9868-0.0132=0.9736 \text{ (from z-table).}"